# Diffusion approximation of the stationary distribution of a two-level single server queue

### v3

@@ 5.1 Process limit
Recall that one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation, that is, BAR:
@@
which follows from (4.3), (4.54) and (4.63) since $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$. For $b_{1}=0$, BAR is obtained from (5.1) by letting $b_{1}\downarrow 0$. So, we only consider the case that $b_{1}=0$.
@@
where $W(\cdot)$ is the standard Brownian motion, $Y(\cdot)\equiv\{Y(t);t\geq 0\}$ is a non-deceasing process satisfying that $\int_{0}^{t}1(Z(u)>0)Y(du)=0$ for $t\geq 0$. If $Y(\cdot)$ is absolutely continuous with respect to Lebesgue measure on $\mathbb{R}_{+}$, then (5.2) has a unique weak solution by Theorem 5.15 of [7] because $\sigma(x)\geq\sqrt{c_{1}}\sigma_{1}\wedge\sqrt{c_{2}}\sigma_{2}>0$.
However, $Y(\cdot)$ is nondecreasing, but may not be absolutely continuous. So far, we just assume that there is a weak solution of (5.2). For notational convenience, we continue to use the same notations for this weak solution. To apply Ito integration rule to (5.2), let us compute the quadratic variation $\big{[}Z\big{]}_{t}$ of $Z(\cdot)$. Define $M(\cdot)\equiv\{M(t);t\geq 0\}$ as
[…]
then $M(\cdot)$ is a martingale. Denote the quadratic variations of $Z(\cdot)$ and $M(\cdot)$, respectively, by $\big{[}Z\big{]}_{t}$ and $\big{[}M\big{]}_{t}$. Then, $Z(t)$ and $Y(t)$ are continuous in $t$, so it follows from (5.2) that
[…]
We next define local time $L_{x}(t)$ for $x\geq 0$ as
[…]
where ${\rm sgn}(x-)=1(x>0)-1(x\leq 0)$ (e.g., see (4) of Chapter 22 of [6]). Note that, for any measurable function $f\geq 0$,
[…]
by Theorem 22.5 of [6]. Note that $L_{x}(t)$ is $a.s.$ continuous in $t\geq 0$, but may not be so in $x$ if $Z(\cdot)$ is not a martingale. So, let $L^{+}_{x}(t)$ and $L^{-}_{x}(t)$ be the right and left continuous versions on $x$, respectively, of $L_{x}(t)$ for each fixed $t>0$. Namely, $L^{+}_{x}(t)=\lim_{\varepsilon\downarrow 0}L_{x+\varepsilon}(t)$ and $L^{-}_{x}(t)=\lim_{\varepsilon\downarrow 0}L_{x-\varepsilon}(t)$. As we will see, their difference is a key for our analysis.
To derive a BAR for (5.2), we use the extended Ito formula for a convex test function $f$ for the local time $L^{-}_{x}(t)$, which is given as
[…]
where $f^{\prime}_{+}(x)=f^{\prime}(x+)$, that is, left derivative of $f$ at $x$, and $\mu_{f^{\prime}_{+}}$ is a measure on $(\mathbb{R}^{+},{\mathcal{B}}(\mathbb{R}_{+}))$ defined by
$\mu_{f^{\prime}_{+}}((x,y])=f^{\prime}_{+}(x)-f^{\prime}_{+}(y),\qquad x<y\mbox{ in }\mathbb{R}_{+}.$
Note that (5.7) is the left-continuous local time counterpart of Theorem 22.5 of [6], which takes the right-continuous local time $L^{+}_{x}(t)$ and $\mu_{f^{\prime}_{-}}$ instead of $L^{-}_{x}(t)$ and $\mu_{f^{\prime}_{+}}$.
For (5.7), we take the test function $f$ with parameter $\theta\geq 0$, which is defined as $f(x)=e^{\theta h^{-}_{\ell_{0}}(x)}$, where $h^{-}_{\ell_{0}}(x)=(\ell_{0}-x)1(0\leq x\leq\ell_{0})$. Then, $f$ is a convex function for $x\geq 0$, and $f^{(n)}_{+}(x)=\theta^{n}e^{\theta x}1(0\leq x<\ell_{0})$ for $n\geq 1$. Since $\mu_{f^{\prime}_{+}}(\{\ell_{0}\})=-f^{\prime}(\ell_{0}-)$ and $\mu_{f^{\prime}_{+}}$ has the density $f^{\prime\prime}_{+}(x)=\theta^{2}e^{\theta x}1(0\leq x<\ell_{0})$, it follows from (5.4) and (5.6) that
[…]
$=\frac{1}{2}\theta^{2}\int_{0}^{t}e^{\theta Z(u)}1(0\leq Z(u)<\ell_{0})d[Z]_{u}-\frac{1}{2}\theta e^{\theta\ell_{0}}L^{-}_{\ell_{0}}(t).$
Hence, the extended Ito formula (5.7) becomes
[…]
If $Z(\cdot)$ is a stationary process, then taking the expectation of (5.1) for $t=1$ yields
$\Big{(}-b_{1}\theta+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta^{2}\Big{)}\widetilde{\varphi}_{1}(\theta)-\frac{1}{2}\theta e^{\theta\ell_{0}}\mathbb{E}[L^{-}_{\ell_{0}}(1)]+\theta\widetilde{\varphi}_{1}(0)\mathbb{E}[Y(1)]=0,$ (5.9)
where $\widetilde{\varphi}_{1}(\theta)$ is the moment generating function of $Z(1)$ on $0\leq Z(1)\leq\ell_{0}$.
Similarly, let $\widetilde{\varphi}_{2}(\theta)$ be the moment generating function of $Z(1)$ on $Z(1)>\ell_{0}$, then, taking the right-continuous version $L^{+}_{x}(t)$ of the local time $L_{x}(t)$, and letting $f(x)=-e^{\theta h^{+}_{\ell_{0}}}$ for $\theta\leq 0$, where $h^{+}_{\ell_{0}}=(x-\ell_{0})1(x>\ell_{0})$, we have
$\Big{(}-b_{2}\theta+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta^{2}\Big{)}\widetilde{\varphi}_{2}(\theta)+\frac{1}{2}\theta e^{\theta\ell_{0}}\mathbb{E}[L^{+}_{\ell_{0}}(1)]=0.$ (5.10)
We finally note that
[…]
Then, combining (5.9) and (5.10), we arrive at (5.1). Thus, the BAR of $Z(\cdot)$ is identical with (5.1). Based on this fact, we conjecture that
@@
    The $\widetilde{\nu}$ of (3.7) in Theorem 3.1 is the unique solution of the BAR (5.1) and the stationary distribution of $Z(t)$.
These conjectures are challenging, but we expect that they are verifiable.
@@
$\mathbb{E}[B^{(r)}_{2}]\sim\frac{\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{c^{(r)}_{2}\mu^{(r)}(1-\rho^{(r)}_{2})}=\frac{c_{2}}{c_{1}+c_{2}}\times\frac{1}{rb_{2}\mu}+o(1).$ (5.13)
@@
which is compatible with (4.59). Hence, (A.9) yield
@@
$=\frac{\beta_{2}(1-e^{-\beta_{1}\ell_{0}})}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)+c_{2}b_{1}}.$
Thus, this and $\widetilde{A}_{1}+\widetilde{A}_{2}=1$ concludes (4.35).
@@
which is compatible with (4.62). Hence, by $\beta_{2}=b_{2}/c_{2}$,
$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\ell^{(r)}_{0}\pi^{(r)}_{0}=\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}}.$
It remains to verify (3.4) and (3.5). This can be done through computing the limit of the distribution function $H(x)$ below as $r\downarrow 0$.
$H(x)\equiv\mathbb{P}[rL^{(r)}\leq x]=\sum_{n=0}^{[x]}\pi^{(r)}_{n},\qquad x\in\mathbb{R}_{+}.$
However, we omit the verification of (3.4) and (3.5) because they are part (i) of Theorem 3.1.

### v4

@@ 5.1 Process limit
Recall that, if one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation, that is, BAR:
@@
which follows from (4.3), (4.54) and (4.63) since $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$. For $b_{1}=0$, BAR is obtained from (5.1) by letting $b_{1}\downarrow 0$. So, we only consider the case that $b_{1}\not=0$.
@@
where $W(\cdot)$ is the standard Brownian motion, $Y(\cdot)\equiv\{Y(t);t\geq 0\}$ is a non-deceasing process satisfying that $\int_{0}^{t}1(Z(u)>0)Y(du)=0$ for $t\geq 0$. If $Y(\cdot)$ is absolutely continuous with respect to Lebesgue measure on $\mathbb{R}_{+}$, then (5.2) has a unique weak solution by Theorem 5.15 of [6] because $\sigma(x)\geq\sqrt{c_{1}}\sigma_{1}\wedge\sqrt{c_{2}}\sigma_{2}>0$. However, $Y(\cdot)$ may not be absolutely continuous.
So far, in the previous versions of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered the stationary distribution using the generalized Ito formula for a convex function. We got a stationary equation similar to (5.1) for the $Z(\cdot)$, but could not completely obtained the stationary distribution although we had partial information which is compatible with Theorem 3.1. So, it is conjectured in the previous versions that
@@
    the stationary distribution of $Z(t)$ uniquely exists under the stability condition $b_{2}>0$, and agrees with $\widetilde{\nu}$ of (3.7) in Theorem 3.1 under the one of conditions in (ii) of it in addition to $b_{2}>0$.
During the reviewing process of this paper, the author have known that (a) is resolved by Atar et al. [1, Lemma 4.3] (see also [12] for a different proof), and (c) is affirmatively answered in [12]. Furthermore, the stationary distribution of $Z(t)$ which is obtained in [12] does not need any of the extra conditions in Theorem 3.1. This is not surprising because limiting operation for $r\downarrow 0$ is not needed to derive the stationary distribution of $Z(t)$. Thus, we need not further discuss about the conjectures (a) and (c). As for the conjecture (b), it is still open (see [12]).
@@
$\mathbb{E}[B^{(r)}_{2}]\sim\frac{\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{c^{(r)}_{2}\mu^{(r)}(1-\rho^{(r)}_{2})}=\frac{c_{2}}{c_{1}+c_{2}}\times\frac{1}{rb_{2}\mu}+o(1),$ (5.3)
where $\sim$” stands for the same order as $r\downarrow 0$, that is, for non-zero valued functions $f_{1}$ and $f_{2}$, $f_{1}(r)\sim f_{2}(r)$ means that $\lim_{r\downarrow 0}f_{1}(r)/f_{2}(r)=1$.
@@
which is compatible with (4.59). It follows from (A.9) and (A.2) that
@@
$=\frac{\beta_{2}(1-e^{-\beta_{1}\ell_{0}})}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}=\frac{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)}{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)+c_{2}b_{1}},$ (A.13)
where the last equality uses the fact that $\beta_{i}=b_{i}/c_{i}$ for $i=1,2$.
@@
which is compatible with (4.62). Since $\beta_{2}=b_{2}/c_{2}$, (A.9) and (A.14) yield
$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\ell^{(r)}_{0}\pi^{(r)}_{0}=\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}}.$ (A.15)
Hence, from (3.14), (A.2) and (A.15) and using the fact that $\nu^{(r)}_{1}(\mathbb{R}_{+})+\nu^{(r)}_{2}(\mathbb{R}_{+})=1$, we have
$\lim_{r\downarrow 0}\nu^{(r)}_{i}(\mathbb{R}_{+})$ $=\widetilde{A}_{\exp,i},\qquad i=1,2.$ (A.16)
It remains to verify (3.4) and (3.5). This can be done through the Laplace transforms of $\{\pi^{(r)}_{n};0\leq n\leq\ell^{(r)}_{0}\}$ and $\{\pi^{(r)}_{n};n>\ell^{(r)}_{0}\}$ together with (A.16). We omit its detail because they are just computations.
@@
Acknowledgement
I thank the editor and referees for their helpful comments and careful reading. This study is motivated by the BAR approach, coined by Jim Dai (e.g., see [3]). I am grateful to him for his grand design about the BAR approach, and to Rami Atar for providing his recent paper [1]. I also benefit from Evsey Morozov for his comments on this paper.